OpenStudy (anonymous):

The height of a ball dropped from a tall building is modeled by the equation d(t) = 16t2 where d equals the distance traveled at time t seconds and t equals the time in seconds. What does the average rate of change of d(t) from t = 2 to t = 5 represent? a.)The ball falls down with an average speed of 336 feet per second from 2 seconds to 5 seconds. b.)The ball falls down with an average speed of 48 feet per second from 2 seconds to 5 seconds. c.)The ball travels an average distance of 336 feet from 2 seconds to 5 seconds. d.)The ball travels an average distance of 48 feet from 2 seconds

OpenStudy (math_genius123):

i think its B.

OpenStudy (anonymous):

why?

OpenStudy (math_genius123):

umm.... ok just a sec

OpenStudy (math_genius123):

its either A or B because its being dropped

OpenStudy (math_genius123):

so its falling

OpenStudy (math_genius123):

umm... so either A or B

OpenStudy (anonymous):

find d 1. when t=2 s 2.when t=5s 3. find difference 4.this gives distance covered in (5-2)=3 s 5.divide by ( 5-2)=3s this gives average speed

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

d(t) = 16 * t^2 when t = 2: d(2) = 16 * 2^2 = ? when t = 5: d(5) = 16 * 5^2 = ? Distance traveled between t = 2 and t = 5 is: d(5) - d(2) = ? Average speed = { d(5) - d(2) } / ( 5 - 2 ) = ?

OpenStudy (anonymous):

d(3)

OpenStudy (ranga):

Just plug the numbers into a calculator and fill out all the question marks in my previous reply.

OpenStudy (anonymous):

d(2) = 16 * 2^2 = 64 when t = 5: d(5) = 16 * 5^2 = 400 Distance traveled between t = 2 and t = 5 is: d(5) - d(2) = 336 Average speed = { d(5) - d(2) } / ( 5 - 2 ) = 112

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

Yes. See which choice fits the answer.

OpenStudy (ranga):

You found the average SPEED to be 112 feet / sec. You found the average DISTANCE traveled to be 336 feet. Which choice is the correct answer?

OpenStudy (ranga):

The first two choices deal with average SPEED. The last two choices deal with average DISTANCE.

OpenStudy (anonymous):

c? @ranga

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

thanks. can you help me with one more?

OpenStudy (ranga):

I can guide you but I cannot give the answer.

OpenStudy (anonymous):

functions 1 and 2 are shown below function 1 f(x)=-3x^2+2

OpenStudy (anonymous):

function 2

OpenStudy (anonymous):

which function has a larger maximum?

OpenStudy (ranga):

The general equation of a parabola is ax^2 + bx + x. The maximum or minimum occurs at x = -b/(2a) For function 1) find at what value of x the maximum occurs by finding -b/(2a) Then put that x value in the function and find the maximum value of the function. What do you get?

OpenStudy (ranga):

It is ax^2 + bx + c (not x)

OpenStudy (anonymous):

im confused :/

OpenStudy (ranga):

Compare ax^2 + bx + c to -3x^2 + 2 what is a, b and c?

OpenStudy (anonymous):

a=-3x^2 b= 1x c=2

OpenStudy (ranga):

a,b,c are coefficients. So a = -3, b = 0 and c = 2 Maximum/minimum for a parabola occurs at x = -b/(2a) plug the values for a and b and find what x value the max/min occurs.

OpenStudy (anonymous):

x=1?

OpenStudy (ranga):

a = -3, b = 0, c = 2 What is -b/(2a)?

OpenStudy (anonymous):

-0/2-3?

OpenStudy (ranga):

-0 / (2 * -3) = -0 / (-6) = ?

OpenStudy (anonymous):

0?

OpenStudy (ranga):

Yes, the maximum occurs at x = 0 Put x = 0 in the first function to find what the maximum value is.

OpenStudy (anonymous):

f(x)=-3x^2+2 f(0)=-3(0)^2+2?

OpenStudy (ranga):

simplify above. what is the max value?

OpenStudy (anonymous):

2

OpenStudy (ranga):

Yes. The max value of function 1 is 2. From the graph, what is the max value of function 2?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

I think

OpenStudy (ranga):

Yes. So which function has the larger maximum?

OpenStudy (anonymous):

function 2?

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

thanks for all your help :)

OpenStudy (ranga):

you are welcome.